Professor Blockhus gives two different statistics tests, but one test is harder than the other. Scores on test A are normally distributed with a mean score of 78 and a standard deviation of 6. Scores on test B are also normally distributed but with a mean score of 65 and a standard deviation of 9.

If Erik scored an 79 on test B, what percent of the class scored below him? That is, what is his percentile on test B.
Answer #1

z score for Erik score in test B = (79 - 65)/9 = 1.56

Hence,

Percent of class scores below him

= P(z < 1.56)

= 0.94

= 94%